A dose of methylnaltrexone bromide is administered to patients weighing over 251 lbs at a level of 0.15 mg per kilogram of body mass. Each dose is administered through a subcutaneous injection every other day. If a patient weighs 264 lbs and the methylnaltrexone bromide solution has a concentration of 20 mg/ml, how many milliliters will be administered in a period of two weeks?
Answer #1

ANSWER.

* Calculation ml of drug required in a period of two weeks.

        Formula

             Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

           Recommended dose = 0.15mg/kg.

           Weight of the patient = 264 lbs.

      Convert lbs to kg for calculation

                     1 lbs = 0.454 kg

        So

            Weight of the patient = 264 * 0.454

                                              = 119.7 kg

             Desired dose (D) for this patient = 0.15 * 119.7

                                                                  = 17.9 mg.

                          Dose in hand (H) = 20 mg

                                              Quantity (Q) = 1ml

         As per above formula

                  Volume(ml) of drug required = D/H * Q

                                                                = 17.9/20 * 1

                                                                = 0.9 ml /ḍay.

       ml of drug required in a period of two weeks or 14 days.

              Dose is administered every other day means first day,one day day off,

              then next day.

                 (Day 1,day3,day5,day7,day9,day11,day13) - total 7 days.

         So

             ml of drug required in a period of 2 weeks = 0.9 * 7

                                                                                       = 6.3 ml.